zniyab1:

A spacecraft is traveling in interplanetary space at a constant velocity. What is the estimated distance traveled by the spacecraft in 9.50 seconds? 2.30 × 102 kilometers 2.42 × 102 kilometers 2.51 × 102 kilometers 3.02 × 102 kilometers

justjm:

What is the magnitude of your velocity? Then just plug it into v=d/t, where d=displacement. You can only find distance here since it is a constant velocity

zniyab1:

?

zniyab1:

so whats tthe answer

justjm:

Are you given a diagram or additional information? Your velocity is arbitrary at the moment because the magnitude nor the direction is given, but more specifically the magnitude.

justjm:

Other than that, we are forced to assume that the spacecraft is following a velocity of 242 m/s. So just plug that in. \(242~m/s=\frac{d}{9.50}\) solve for d